A two digit number is 6 times its units digit. Find the number if the sum of its digits is 6.
Answer by ankor@dixie-net.com(22740)

Let x = the 10's digit
Let y = the units
(10x + y) is the two digit number
:
Write an equation for each phrase
"A two digit number is 6 times its units digit."
10x + y = 6y
10x = 6y - y
10x = 5y
simplify, divide by 5
2x = y
:
"the sum of its digits is 6."
x + y = 6
replace y with 2x
x + 2x = 6
3x = 6
x = 2
therefore
y = 4
:
24 is the number
:
:
Check 24 = 6*4